One number is three times a second number, and twice the sum of the numbers is 168. Find the Numbers.
Answer by mananth(16949)

second number = x
first number = 3x
2(3x+x)= 168
/2
4x=84
/4
x=21
other number = 3x = 3*21= 63
..
21, 63
..
m.ananth@hotmail.ca